For a given body at a particular temperature, its coefficients of reflection and transmission are 0.63 and 0.21 respectively. Then its coefficient of emission at the same temperature will be
(a) 0.16
(b) 0.32
(c) 0.84
(d) 0.36

Grade:Upto college level

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

0.63+0.21+x= 1

x=0.16
option a

Regards
Vikas TU
3 years ago
Given :   r=0.3     a=0.2    
Using   a+r+t=1
So,   0.2+0.3+t=1
⟹ t=0.5
…...............................................................